5x1 -4x2 +3x3 +3x4 = 0
-x1 +x2 +2x3 +2x4 = 2
4x1 -3x2 +5x3 +5x4 = 2
-3x1 +3x2 +6x3 +6x4 =6
x1
x2=(results)+s(results)+t(results)
x3
x4

Answer by greenestamps(13200)



5x1 -4x2 +3x3 +3x4 = 0
-x1 +x2 +2x3 +2x4 = 2
4x1 -3x2 +5x3 +5x4 = 2
-3x1 +3x2 +6x3 +6x4 =6

(1) Use a calculator like a TI-83 that can work with matrices.

(2) Use Cramer's Rule.

(3) If you want to do it algebraically, the process is simple, but tedious. Use elimination to reduce the system of 4 equations and 4 unknowns to a system of 3 equations in 3 unknowns, then repeat the process to get 2 equations in 2 unknowns, and finally solve that pair of equations.

There are innumerable different paths you could take to get the answer that way; our taking the time to show you a particular path would not teach you anything useful.

If you have a particular question about how to solve a problem like this, ask it; and show or at least describe any work you have done or what troubles you are having. We aren't going to do your work for you if you just show the system of equations and wait for us to do it.

------------------------------------------------------------

To the student....

Re-post the problem and specify that you are continuing a discussion with me so that you want me (greenestamps) to be the one to respond.

In your post, show the work you did to get your values for x1, x2,x3, and x4 (they aren't right). Also explain the "+s and +t" and the "other columns" you are talking about. I assume it has something to do with how to display your answers....